OpenStudy (anonymous):

Write an equation of a circle if the endpoints of the diameter of the circle are (0,3) and (0,-4)

OpenStudy (p0sitr0n):

d= -4 - 3 = -7 circle equation (x-h)^2 + (y-k)^2 = r^2 x^2 + y^2 = 3.5^2

OpenStudy (anonymous):

I don't get it cuz I got (x-0)^2+(y-1/2)^2=2.5

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

So how do u do it

OpenStudy (anonymous):

let me work on it one sec

OpenStudy (anonymous):

x^2+(y+0.5)^2=(7/2)^2

OpenStudy (anonymous):

ya, that is correct

OpenStudy (anonymous):

how did you get that though?

OpenStudy (anonymous):

Rotate it around to the x-axis, apply the formula and swap x and y back again.

OpenStudy (anonymous):

on scratch paper?

OpenStudy (anonymous):

Sure, try it, it's easier to do than explain...

OpenStudy (anonymous):

and what formula?

OpenStudy (anonymous):

(x-a)^2+(y-b)^2 = r^2

OpenStudy (anonymous):

oh, sorry, i didn't read that up there

OpenStudy (anonymous):

ok, thank you for your explanation, it has been very enlightening.

OpenStudy (anonymous):

u r welcome

OpenStudy (anonymous):

Thanks to both of u